OpenStudy (anonymous):

(15/x^2+3x)+(3/x)+(5/x+3) Simplify if possible

OpenStudy (anonymous):

first, find the common denominator and multiply each term by the common denominator to rid off the fraction, then you can easily simplify it

OpenStudy (anonymous):

I got \[(8x^3+48x^2+74x)/((x+3)(x^3+3x^2))\]

OpenStudy (anonymous):

(x^2+14x+18)/(x^2+3x)

OpenStudy (anonymous):

how did you arrive to that answer

OpenStudy (anonymous):

right or wrong

OpenStudy (mertsj):

\[\frac{15}{x(x+3)}+\frac{3}{x}+\frac{5}{x+3}\]

OpenStudy (mertsj):

\[\frac{15}{x(x+3)}+\frac{3(x+3)}{x(x+3)}+\frac{5x}{x(x+3)}\]

OpenStudy (mertsj):

Now, all the denominators are the same so we will simply add the numerators:

OpenStudy (mertsj):

\[\frac{15+3(x+3)+5x}{x(x+3)}=\frac{15+3x+9+5x}{x(x+3)}=\frac{8x+24}{x(x+3)}\]

OpenStudy (anonymous):

Thank you so much I greatly appreciate your help, you're the best.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

good Mertsj

OpenStudy (mertsj):

Whoops. Not quite done: \[\frac{8x+24}{x(x+3)}=\frac{8(x+3)}{x(x+3)}=\frac{8}{x}\]

OpenStudy (anonymous):

Good looking out =^^=